OpenStudy (anonymous):

Explain the difference between using the trigonometric ratios (sin, cos, tan) to solve for a missing angle in a right triangle versus using the reciprocal ratios (sec, csc, cot). You must use complete sentences and any evidence needed (such as an example) to prove your point of view.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Do you know what the reciprocal ratios are?

OpenStudy (whpalmer4):

I've never seen a catchy mnemonic for them, but \[\cot x = \frac{1}{\tan x}\]\[\sec x = \frac{1}{\cos x}\]\[\csc x = \frac{1}{\sin x}\]

OpenStudy (whpalmer4):

That means that \[\cot x = \frac{\text{Adjacent}}{\text{Opposite}}\]\[\sec x = \frac{\text{Hypotenuse}}{\text{Adjacent}}\]\[\csc x = \frac{\text{Hypotenuse}}{\text{Opposite}}\]

OpenStudy (anonymous):

will this be my answer??

OpenStudy (whpalmer4):

I'm just explaining the idea to you, not writing your answer. So if we have that same triangle, and we wanted to find \(\csc x\), what would it be?

OpenStudy (anonymous):

hypostuse/oppisote

OpenStudy (whpalmer4):

okay, but what would the value be?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

csx

OpenStudy (anonymous):

?

OpenStudy (whpalmer4):

What is the hypotenuse in that triangle? What is the opposite side from angle \(x\)?

OpenStudy (whpalmer4):

\[\csc x = \frac{\text{Hypotenuse}}{\text{Opposite}}\]

OpenStudy (anonymous):

thhe adjacent side

OpenStudy (whpalmer4):

numbers, man, numbers! Look at the triangle?